A 3.00-kg block is placed on top of a 12.0-kg block that rests on a frictionless table. The coefficient of static friction between the two blocks is0.500. What is the maximum horizontal force that can be applied before the 3.00-kg block begins to slip relative to the 12.0-kg block, if the force is applied to each of the following?

(a) Find the acceleration with which Robin is pulled upward.

(b) Find the tension in the rope while Robin escapes.
Answer #1

Maximum horizontal force= Mg*mu

=3*10*0.5

=15 N

Accelation=

15 N=m*a

15/12=a=1.25 m/s^2

Tension T=Mg+ma

=12*10+3*1..25

=123.75 N